Circular No. 6370

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 B2 (HYAKUTAKE) P. D. Feldman, Johns Hopkins University; M. Festou, Observatoire de

Midi-Pyrenees, for the IUE ToO Team for Comet Observations; and P. M. Rodriguez and R. Gonzalez, European Space Agency IUE Observatory, report: “Further observations (IAUC 6355) were made in both the SWP and LWP range on Mar. 27.6 and 27.7 UT with the International Ultraviolet Explorer (IUE) on the nuclear region of comet C/1996 B2. The observations clearly show the presence of the strongest CO Cameron bands at 199.3, 206.3, and 215.8 nm, as well as the CO^{+} first negative bands at 211.4 and 219.1 nm. At this time, the water-production rate was derived as 2.4 × 10^{29} s^{−1}, which is close to the level seen on Mar. 23.12, and significantly less than observed on Mar. 25.06, suggesting that the nucleus lost some small pieces. The Cameron bands are produced primarily by the photodissociation of CO_{2} and a preliminary estimate of Q(CO_{2}) is 2.5 × 10^{28} s^{−1}. The CO-production rate is found to be 7 × 10^{27} s^{−1}, derived from the observed optically-thin fourth positive bands at 154.5 and 159.7 nm. This is considerably lower than the CO-production rate derived from sub-millimeter observations earlier in the month (IAUC 6353), and suggests that most of the CO in the coma is derived from a distributed source, as was the case for 1P/Halley.”

Visual m_{1} estimates, made in strong moonlight: Apr. 1.21 UT, 2.1 (C. E. Spratt, Victoria, BC, naked-eye); 1.84, 1.8 (M. L. Paradowski, Lublin, Poland, 7×35 binoculars); 2.06, 2.0 (G. W. Kronk, Troy, IL, naked eye).

COMET 29P/SCHWASSMANN-WACHMANN 1 I. Ferrin, Universidad de los Andes, Merida, reports that CCD images

of comet 29P were taken in V, R, and I on Mar. 24.35 UT with the 1m Schmidt telescope of the National Observatory of Venezuela, showing a coma diameter of 216′′ and an intense jet between p.a. 238^{o} and 288^{o} (maximum intensity at p.a. 269^{o}; solar direction 281^{o}). There is also the remains of a second jet at p.a. 338^{o}. Preliminary photometry of the whole coma gives V = 12.9 ± 0.15.

Further total visual magnitude estimates (cf. IAUC 6351): Mar. 24.05 UT, 11.1 (J. Carvajal, Teruel, Spain, 0.44-m reflector); 27.88, 10.9 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector).

1996 April 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green